# Scrutinize What We Ignore: Reining Task Representation Shift In Context-Based Offline Meta Reinforcement Learning

### v1

@@ Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that maximizing the mutual information between the task and the task representation ($I(Z;M)$) can lead to performance improvements. Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking. Motivated by the return discrepancy scheme in the model-based RL field, we find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. However, this optimization process ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we turn to use the framework of performance difference bound to consider the impacts of task representation shift explicitly. We demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against previous approaches. To make it practical, we design an easy yet highly effective algorithm RETRO (REining Task Representation shift in context-based Offline meta reinforcement learning) with only adding one line of code compared to the backbone. Empirical results validate its state-of-the-art (SOTA) asymptotic performance, training stability and training-time consumption on MuJoCo and MetaWorld benchmarks.
@@ 1 Introduction
Reinforcement Learning (RL) has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering], primarily through online RL methods. However, the extensive interactions of online RL entail high costs and safety concerns, posing significant challenges in real-world applications. Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar]. Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
@@
As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting, li2024towards] has emerged as a beacon to illuminate the way toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Along the line of OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] wield a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, they propose to train a context encoder independently to learn the task representation from the collected context (e.g. trajectories), then pass the latent representation to the downstream policy and value function, forming a decoupled framework. Thanks to [li2024towards] for unveiling the nature of the prior COMRL works that they predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements [li2020focal, yuan2022robust, gao2024context, li2024towards]. Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking.
Intriguingly, we unify the mutual information maximization objective commonly used in COMRL with the general RL objective of maximizing expected return. To be specific, maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee, filling the gap between intuition and theoretical justification. However, after scrutinizing this optimization process, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we adopt the framework of performance difference bound to consider the impacts of task representation shift explicitly. We theoretically demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against prior works.
To make it practical, we derive an optimization problem concerning task representation shift and find that reining the update frequency of task representation and employing a straightforward classifier-based method for task representation extraction is sufficient to beat existing SOTA algorithms. RETRO is evaluated on a wide range of RL benchmarks covering MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta]. With the addition of just one line of code, RETRO achieves superior asymptotic performance and training stability compared to the classifier-based method alone. Notably, RETRO also holds a tuning-free property. Moreover, our approach to reining task representation shift is general and is also ablated in the experiments.
@@ 2 Related Works
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of learned task representation to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], many prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements. Though achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking. We introduce the return discrepancy scheme [janner2019trust] into the COMRL setting to provide a feasible explanation. However, after scrutinizing this optimization process, we further find that it ignores the impacts of task representation shift, which may lead to performance improvement collapse. Centered around this motivation, our work aims to present a theoretical framework that reins the task representation shift.
Performance Improvement Guarantee. Performance improvement guarantee is a fundamental concern in both online RL and offline RL settings. In online RL, performance difference bound [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic] and regret bound [osband2014model, curi2020efficient] are widely adopted schemes to get a performance improvement guarantee. In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the multi-task and COMRL settings largely unexplored. We first provide a theoretical justification w.r.t the performance improvement guarantee for prior COMRL works. Then, we prove that it is possible to achieve monotonic performance improvements by reining the task representation shift ignored by prior COMRL works, thereby showcasing the advantage against them.
@@ 3.1 Problem Statement
A task in RL is formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$. A viable solution is to iterate the Bellman operator over the action-value function [sutton2018reinforcement], which is denoted as $Q_{\pi}(s,a)=\mathbb{E}_{\pi,P}[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
It is widely assumed that the task $m$ in the meta RL setting is subject to the distribution $p(M)$ with sharing the same state space and action space but different from transition dynamics and reward functions [li2020focal, yuan2022robust, gao2024context, li2024towards]. At training time, tasks are randomly sampled from the task distribution $p(M)$. We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with the only right to access the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
@@ 3.2 Context-based OMRL
COMRL holds a steadfast belief that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, COMRL methods would train a task encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exist the optimal task representation distribution $Z^{*}(\cdot|x)$ that can deduce the optimal expected return. The objective of COMRL is formalized as:
@@
Prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvement, as shown in Theorem 3.1.
@@
Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. When meeting $I(Z;M;X_{b})\geq 0$:
@@
However, the theoretical justification of performance improvement for such intuition has been lacking.
@@ 4 Methods
To fill the gap, we introduce the return discrepancy scheme [janner2019trust] into the COMRL setting and find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation (Section 4.1). Further, after scrutinizing this optimization process, we argue that it ignores the impacts of task representation shift, which may lead to performance improvement collapse. We then turn to use the framework of performance difference bound to consider task representation shift explicitly. We theoretically prove that it is possible to achieve monotonic performance improvements by reining task representation shift, thereby showcasing the advantage against prior works (Section 4.2). To make it practical, we derive an optimization problem concerning task representation shift and use the simplest algorithm to instantiate it (Section 4.3).
We first introduce the Lipschitz assumption for the following use. Lipschitz-based assumptions are widely adopted in both model-free RL [song2019efficient, ghosh2022provably] and model-based RL [ji2022update, zhang2023how].
@@
The policy function is $L_{z}-$ Lipschitz with respect to some norm |||||$\cdot$||| | ⋅ | | in the sense that
@@ 4.1 A Performance Improvement Perspective Towards Prior Works
From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the task encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage iteration framework with the first stage optimizing the task encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage optimizing the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, which forms a bedrock to connect the bridge between the intuition of prior works and the theoretical justification concerning the performance improvement guarantee.
@@
Theorem 4.3 and Eq. (6) indicate the following two principles for the task encoder and the policy. For the task encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation is irrelevant for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
@@
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the task representation shift between two consecutive updates alone. In the next section, we will show that the impacts of task representation shift may lead to performance improvement collapse.

### v2

# Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning

@@ Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task and the task representation ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored. Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby providing a feasible explanation concerning performance improvements. Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating optimization process, which may lead to performance improvement collapse. We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates. We set different manners to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
@@ 1 Introduction
RL has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering, zhang2023safe], primarily through online RL methods. However, the extensive interactions required by online RL entail high costs and safety concerns, posing significant challenges for real-world applications. Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar]. Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
@@
As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting] has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Among the OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, these methods [li2020focal, gao2024context, li2024towards] propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g. trajectories) and train the downstream policy with standard offline RL algorithms conditioning on the task representation, as shown in Figure 1 left. The context encoder and policy are trained in an alternating manner, with each being updated once per cycle. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
Figure 1: Our training framework compared to the previous training framework. They both adopt the alternating optimization framework to train the context encoder and the policy. However, ours considers the variation of task representation ignored by the previous training framework.
Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return. To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks. This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process. Variations of the task representation that are either too large or too small can lead to performance improvement collapse. To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement guarantees with proper updates.
To show the impacts of this issue, we set different manners to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities. These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards]. We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
@@ 2 Related Works
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$. However, our work moves the focus on how to get better asymptotic performance based on a given context encoder learning algorithm. Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
Performance Improvement Guarantee.  Ensuring performance improvement is a key concern in both online and offline reinforcement learning settings. In online RL, performance improvement guarantees are often established through methods such as performance difference bounds [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic], and regret bounds [osband2014model, curi2020efficient]. In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the COMRL settings largely unexplored. While ContrBAR [choshen2023contrabar] also benefits from the performance improvement guarantee in the OMRL setting, it is specifically served for its own framework. In contrast, our work addresses a broader class of algorithms that maximize $I(Z;M)$, including but not limited to ContrBAR.
@@ 3.1 Problem Statement
A task in RL is generally formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$.
Given a distribution of tasks $p(M)$, every task $m$ sampled from $p(M)$ is an MDP sharing the same state space and action space but differing from transition dynamics and reward functions. It is widely assumed that the tasks in the meta RL setting are randomly sampled from $p(M)$ [li2020focal, yuan2022robust, gao2024context, li2024towards]. We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with access only to the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
@@ 3.2 Context-based OMRL
Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$. Hence, the objective of COMRL is formalized as:
@@
Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1. They hold the intuition that using such kind of manner to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements. However, the theoretical justification of performance improvement for such intuition has been less explored.
@@
Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. We have:
@@ 4 Methods
We first introduce a Lipschitz assumption, a widely used technique in both model-free [song2019efficient, ghosh2022provably] and model-based [ji2022update, zhang2023how] RL frameworks.
@@
The policy function is $L_{z}-$ Lipschitz w.r.t some norm |||||$\cdot$||| | ⋅ | | in the sense that
@@ 4.1 A Performance Improvement Perspective Towards Prior Works
From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the context encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, connecting the bridge between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
@@
Denote $Z_{mutual}(\cdot|x)$ as the optimal task representation distribution served for maximizing $I(Z;M)$  [^1] Whether the solution of maximizing $I(Z;M)$ is equivalent to the solution of maximizing the expected return is still an open problem. Hence, we introduce an extra notation here.. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
@@
Theorem 4.3 and Eq. (6) indicate the following two principles for the context encoder and the policy. For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation related terms are constants for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
@@
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the variation of the task representation alone. In the next section, we will show that this may lead to performance improvement collapse.